Grade 10Trigonometry

If cosecA = x + 1/4x ; then the value of cosecA + cot A =?

Neha Naik
7 Years agoGrade 10

1 Answer

Arun
7 Years ago
Cosecθ=x+1/4x
∴, cosec²θ=(x+1/4x)²
=x²+2.x.1/4x+1/16x²
=x²+1/2+1/16x²
Now, cosec²θ-cot²θ=1
or, cot²θ=cosec²θ-1
or, cot²θ=x²+1/2+1/16x²-1
or, cot²θ=x²+1/16x²-1/2
or, cot²θ=x²-2.x.1/4x+1/16x²
or, cot²θ=(x-1/4x)²
or, cotθ=+-(x-1/4x)
∴, either, cosecθ+tanθ
=x+1/4x+x-1/4x [when cotθ=x+1/4x]
=2x  
or, cosecθ+cotθ
=x+1/4x-x+1/4x [when cotθ=-(x+1/4x)]
=1/4x+1/4x
=2/4x
=1/2x
 
Regards
Arun